Two angles are contemporary. The sum of the measure of the first angle and the half of the second angle is 82 degrees. Find the measures of the angles.
What is the measure of the smaller angle?
What is the measure of the other angle?

Answer by unlockmath(1688)

Hello,
Let's set up x and y are the 2 angles. Now we can set up two equations as:
x+y=90
x+1/2(y)=82
Multiply the first equation by -1 to get:
-x-y=-90 and add this to the second equation to get:
-1/2(y)=-8 Multiply both sides by -2 to get:
y=16 degrees
Therefore x will be 74 degrees.
Make sense?
RJ
www.math-unlock.com